# R-matrix valued Lax pair for elliptic Calogero-Inozemtsev system

### v2

@@ 8 Appendix: elliptic functions
$\begin{array}[]{l}\displaystyle{\phi(z,u)=\frac{\vartheta^{\prime}(0)\vartheta(z+u)}{\vartheta(z)\vartheta(u)}=\phi(u,z)\,,\quad\mathop{\hbox{Res}}\limits\limits_{z=0}\phi(z,u)=1\,,\quad\phi(-z,-u)=-\phi(z,u)\,,}\end{array}$ (A.1)
@@
$\begin{array}[]{c}\displaystyle{\theta{\left[\begin{array}[]{c}a\\ b\end{array}\right]}(z|\,\tau)=\sum_{j\in{\zeta}}\exp\left(2\pi\imath(j+a)^{2}\frac{\tau}{2}+2\pi\imath(j+a)(z+b)\right)\,,\quad{\rm Im}(\tau)>0\,.}\end{array}$ (A.3)
@@
$\begin{array}[]{c}\displaystyle{\theta_{1}(u|\tau)=\vartheta(u,\tau)=-i\sum_{k\in{\zeta}}(-1)^{k}q^{(k+\frac{1}{2})^{2}}e^{\pi i(2k+1)u},\qquad\theta_{2}(u|\tau)=\sum_{k\in{\zeta}}q^{(k+\frac{1}{2})^{2}}e^{\pi i(2k+1)u},}\\ \\ \displaystyle{\theta_{3}(u|\tau)=\sum_{k\in{\zeta}}q^{k^{2}}e^{2\pi iku},\qquad\theta_{4}(u|\tau)=\sum_{k\in{\zeta}}(-1)^{k}q^{k^{2}}e^{2\pi iku},}\end{array}$ (A.4)
@@
$\begin{array}[]{c}\theta_{2}(u|\tau)=\theta_{1}(u+\frac{1}{2}|\tau),\!\quad\!\!\theta_{3}(u|\tau)=q^{\frac{1}{4}}e^{\pi iu}\theta_{1}(u+\frac{\tau+1}{2}|\tau),\!\quad\!\!\theta_{4}(u|\tau)=-iq^{\frac{1}{4}}e^{\pi iu}\theta_{1}(u+\frac{\tau}{2}|\tau).\end{array}$ (A.5)
@@
$\begin{array}[]{c}\displaystyle{f(z,u)=\phi(z,u)(E_{1}(z+u)-E_{1}(u)),\qquad f(-z,-u)=f(z,u)\,,}\end{array}$ (A.7)
@@
$\begin{array}[]{c}\displaystyle{E_{1}(z)=\frac{\vartheta^{\prime}(z)}{\vartheta(z)}=\zeta(z)+\frac{z}{3}\frac{\vartheta^{\prime\prime\prime}(0)}{\vartheta^{\prime}(0)}\,,\quad E_{2}(z)=-\partial_{z}E_{1}(z)=\wp(z)-\frac{\vartheta^{\prime\prime\prime}(0)}{3\vartheta^{\prime}(0)}\,,}\end{array}$ (A.8)
$\begin{array}[]{c}\displaystyle{E_{1}(-z)=-E_{1}(z)\,,\quad E_{2}(-z)=E_{2}(z)\,,}\end{array}$ (A.9)
@@
$\begin{array}[]{c}\displaystyle{\omega_{0}=0\,,\quad\omega_{1}=\frac{1}{2}\,,\quad\omega_{2}=\frac{1+\tau}{2}\,,\quad\omega_{3}=\frac{\tau}{2}}\end{array}$ (A.18)
@@
$\begin{array}[]{c}\displaystyle{\varphi_{0}(z,\hbar)=\phi(z,\hbar)\,,\qquad\varphi_{1}(z,\hbar+\omega_{1})=\phi(z,\omega_{1}+\hbar)\,,}\\ \\ \displaystyle{\varphi_{2}(z,\hbar+\omega_{2})=e^{\pi\imath z}\phi(z,\omega_{2}+\hbar)\,,\qquad\varphi_{3}(z,\hbar+\omega_{3})=e^{\pi\imath z}\phi(z,\omega_{3}+\hbar)\,.}\end{array}$ (A.19)
@@
$\begin{array}[]{c}\displaystyle{\varphi_{k}(z,\hbar+\omega_{k})=\frac{\theta_{1}^{\prime}(0)\theta_{k+1}(z+\hbar)}{\theta_{1}(z)\theta_{k+1}(\hbar)}\,,\quad k=0,1,2,3}\end{array}$ (A.20)
The set of introduced functions is transformed under the action of matrix $I$ (3.18)-(A.22) as
@@
$\begin{array}[]{c}\displaystyle{I_{km}=\frac{1}{2}\exp\Big{(}4\pi\imath\Big{(}\omega_{m-1}\partial_{\tau}\omega_{k-1}-\omega_{k-1}\partial_{\tau}\omega_{m-1}\Big{)}\Big{)}\,,\quad k,m=1,...,4\,.}\end{array}$ (A.22)
@@
$\begin{array}[]{l}\displaystyle{\phi(z+1,u)=\phi(z,u)\,,\qquad\phi(z+\tau,u)=e^{-2\pi\imath u}\phi(z,u)}\end{array}$ (A.25)
@@
$\begin{array}[]{l}\displaystyle{E_{2}(z+2\omega_{a})=E_{2}(z)\,,\qquad E_{1}(z+2\omega_{a})=E_{1}(z)-4\pi\imath\partial_{\tau}\omega_{a}}\end{array}$ (A.29)
@@
$\begin{array}[]{l}\displaystyle{E_{1}(\omega_{a})=-2\pi\imath\partial_{\tau}\omega_{a}\,,\quad a\neq 0\,.}\end{array}$ (A.30)

### v3

@@ 8 Appendix: elliptic functions
$\begin{array}[]{l}\displaystyle{\phi(z,u)=\frac{\vartheta^{\prime}(0)\vartheta(z+u)}{\vartheta(z)\vartheta(u)}=\phi(u,z)\,,\qquad\mathop{\hbox{Res}}\limits\limits_{z=0}\phi(z,u)=1\,,\qquad\phi(-z,-u)=-\phi(z,u)\,,}\end{array}$ (A.1)
@@
$\begin{array}[]{c}\displaystyle{\theta{\left[\begin{array}[]{c}a\\ b\end{array}\right]}(z|\,\tau)=\sum_{j\in\mathbb{Z}}\exp\left(2\pi\imath(j+a)^{2}\frac{\tau}{2}+2\pi\imath(j+a)(z+b)\right)\,,\qquad{\rm Im}(\tau)>0\,.}\end{array}$ (A.3)
@@
$\begin{array}[]{c}\displaystyle{\theta_{1}(u|\tau)=\vartheta(u,\tau)=-i\sum_{k\in\mathbb{Z}}(-1)^{k}q^{(k+\frac{1}{2})^{2}}e^{\pi i(2k+1)u},\hskip 18.49988pt\theta_{2}(u|\tau)=\sum_{k\in\mathbb{Z}}q^{(k+\frac{1}{2})^{2}}e^{\pi i(2k+1)u},}\\ \\ \displaystyle{\theta_{3}(u|\tau)=\sum_{k\in\mathbb{Z}}q^{k^{2}}e^{2\pi iku},\hskip 18.49988pt\theta_{4}(u|\tau)=\sum_{k\in\mathbb{Z}}(-1)^{k}q^{k^{2}}e^{2\pi iku},}\end{array}$ (A.4)
@@
$\begin{array}[]{c}\theta_{2}(u|\tau)=\theta_{1}(u+\frac{1}{2}|\tau),\!\qquad\!\!\theta_{3}(u|\tau)=q^{\frac{1}{4}}e^{\pi iu}\theta_{1}(u+\frac{\tau+1}{2}|\tau),\!\qquad\!\!\theta_{4}(u|\tau)=-iq^{\frac{1}{4}}e^{\pi iu}\theta_{1}(u+\frac{\tau}{2}|\tau).\end{array}$ (A.5)
@@
$\begin{array}[]{c}\displaystyle{f(z,u)=\phi(z,u)(E_{1}(z+u)-E_{1}(u)),\hskip 18.49988ptf(-z,-u)=f(z,u)\,,}\end{array}$ (A.7)
@@
$\begin{array}[]{c}\displaystyle{E_{1}(z)=\frac{\vartheta^{\prime}(z)}{\vartheta(z)}=\zeta(z)+\frac{z}{3}\frac{\vartheta^{\prime\prime\prime}(0)}{\vartheta^{\prime}(0)}\,,\qquad E_{2}(z)=-\partial_{z}E_{1}(z)=\wp(z)-\frac{\vartheta^{\prime\prime\prime}(0)}{3\vartheta^{\prime}(0)}\,,}\end{array}$ (A.8)
$\begin{array}[]{c}\displaystyle{E_{1}(-z)=-E_{1}(z)\,,\qquad E_{2}(-z)=E_{2}(z)\,,}\end{array}$ (A.9)
@@
$\begin{array}[]{c}\displaystyle{\omega_{0}=0\,,\qquad\omega_{1}=\frac{1}{2}\,,\qquad\omega_{2}=\frac{1+\tau}{2}\,,\qquad\omega_{3}=\frac{\tau}{2}}\end{array}$ (A.18)
@@
$\begin{array}[]{c}\displaystyle{\varphi_{0}(z,\hbar)=\phi(z,\hbar)\,,\hskip 18.49988pt\varphi_{1}(z,\hbar+\omega_{1})=\phi(z,\omega_{1}+\hbar)\,,}\\ \\ \displaystyle{\varphi_{2}(z,\hbar+\omega_{2})=e^{\pi\imath z}\phi(z,\omega_{2}+\hbar)\,,\hskip 18.49988pt\varphi_{3}(z,\hbar+\omega_{3})=e^{\pi\imath z}\phi(z,\omega_{3}+\hbar)\,.}\end{array}$ (A.19)
@@
$\begin{array}[]{c}\displaystyle{\varphi_{k}(z,\hbar+\omega_{k})=\frac{\theta_{1}^{\prime}(0)\theta_{k+1}(z+\hbar)}{\theta_{1}(z)\theta_{k+1}(\hbar)}\,,\qquad k=0,1,2,3}\end{array}$ (A.20)
The set of introduced functions is transformed under the action of matrix entering (3.18) as
@@
$\begin{array}[]{c}\displaystyle{I_{km}=\frac{1}{2}\exp\Big(4\pi\imath\Big(\omega_{m-1}\partial_{\tau}\omega_{k-1}-\omega_{k-1}\partial_{\tau}\omega_{m-1}\Big)\Big)\,,\qquad k,m=1,...,4\,.}\end{array}$ (A.22)
@@
$\begin{array}[]{l}\displaystyle{\phi(z+1,u)=\phi(z,u)\,,\hskip 18.49988pt\phi(z+\tau,u)=e^{-2\pi\imath u}\phi(z,u)}\end{array}$ (A.25)
@@
$\begin{array}[]{l}\displaystyle{E_{2}(z+2\omega_{a})=E_{2}(z)\,,\hskip 18.49988ptE_{1}(z+2\omega_{a})=E_{1}(z)-4\pi\imath\partial_{\tau}\omega_{a}}\end{array}$ (A.29)
@@
$\begin{array}[]{l}\displaystyle{E_{1}(\omega_{a})=-2\pi\imath\partial_{\tau}\omega_{a}\,,\qquad a\neq 0\,.}\end{array}$ (A.30)
@@ Acknowledgments
No conflict of interest exists for all participating authors. The manuscript has no associated data.